If the measure of X equals 41°, then what is the measure of Y?
Found 2 solutions by Alan3354, Rein:
Answer by Alan3354(69443)

If the measure of X equals 41°, then what is the measure of Y?
-------------
idk
Is Y related to X somehow?

Answer by Rein(1)

it depends.. if the angle is asking for its complement, then the formula will be
y=90-x
y=90-41
y=49
if the angle is asking for its supplement the the formula is
y=180-x
y=180-41
y=139